Consumption = 100+0.8Yd



Investment = 60



Taxes = -150+0.25Y



Government Spending = 80



Find Income and government deficit

1
Expert's answer
2021-11-11T13:25:54-0500

Solution:

Income (Y) = C + I + G

C = 100 + 0.8Yd = 100 + 0.8(Y – T) = 100 + 0.8(Y – (-150 + 0.25Y)) = 220 + 0.6Y

Y = 220 + 0.6Y + 60 + 80

Y – 0.6Y = 220 + 60 + 80

0.4Y = 360

Y = 900

Income (Y) = 900

 

Government deficit = T – G

Government deficit = (-150 + 0.25(900)) – 80

Government deficit = 75 – 80 = -5

Government deficit = -5